Assume: Demand Curve: QD = 80 – 10P; and Supply Curve: QS = 10P 7. Given the information derived above, identify on the graph consumer surplus and producer surplus for each situation as well as deadweight loss, if any. d. Government imposes a sales tax of $4.00, which is equally split between consumers and producers. What is the total surplus? Is the market operating efficiently? INCLUDE A GRAPH
Answer #1

7. Consumer surplus = area between the demand curve and price line = shown by yellow shaded region. Producer surplus = area between the supply curve and price line = shown by green shaded region. There is no deadweight loss before tax.

Demand and supply Consumer Surplus Price Producer Surplus 0 5 10 15 20 25 30 55 60 65 70 75 80 35 40 45 50 Quantity Demand Su

d. Government imposes a sales tax of $4.00, which is equally split between consumers and producers. Now the equilibrium price has risen to $6 and quantity has reduced to 20 units. Consumer surplus and producer surplus are reduced now. There is a government revenue from tax and there is a deadweight loss as well.

Demand and supply Government Price DWE Revenue PS 0 5 10 15 20 55 60 65 70 75 80 25 30 35 40 45 50 Quantity Demand Supply Sup

After tax supply becomes Q = 10(P - 4) or Q = 10P - 40

New equilibrium has

10P - 40 = 80 - 10P

P = 120/20 = $6 and Q = 80 - 60 = 20 units

Producers receive $6 - $4 = $2 as the after tax price

Total surplus = 0.5*(8 - 6)*20 + 0.5*(2 - 0)*20 + (6 - 2)*20 = $120

Market is not efficient since there is a deadweight loss now = 0.5*4*(40 - 20) = $40.